OpenStudy (jjuden):

Solve the inequality |4x-10|>14 . Write the solution in interval notation and graph the solution set on the number line. Show all your work

OpenStudy (jjuden):

\[|4x-10|\ge14\]

OpenStudy (just_one_last_goodbye):

need help?

OpenStudy (jjuden):

yes

OpenStudy (just_one_last_goodbye):

give me a sec ok? ^-^

OpenStudy (just_one_last_goodbye):

Ok done ready?

OpenStudy (jjuden):

yes

OpenStudy (just_one_last_goodbye):

First we add 10 to both sides

OpenStudy (jjuden):

ok

OpenStudy (just_one_last_goodbye):

Btw this is Open Study :) so I need you to help me so I dont get in trouble for giving you a direct answer

OpenStudy (jjuden):

ik

OpenStudy (just_one_last_goodbye):

if we add 10 to both sides what will the equation look like now?

OpenStudy (jjuden):

\[|4x|\ge24\]

OpenStudy (just_one_last_goodbye):

Well it would be a -14

OpenStudy (just_one_last_goodbye):

so... -14 + 10?

OpenStudy (jjuden):

its not -14

OpenStudy (jjuden):

its just 14

OpenStudy (jjuden):

heres the original equation \[|4x-10|\ge14\]

OpenStudy (jjuden):

so if we added 10 to each side it would be \[|4x|\ge24\]

OpenStudy (just_one_last_goodbye):

\[\left| f(x) \right|\ge a \rightarrow f(x) \le -a \]

OpenStudy (jjuden):

ok